OpenStudy (anonymous):

A motorboat travels 248 mi in 4 hours going upstream. It travels 368mi going downstream in the same amount of time. What is the rate of the boat in still water and what is the rate of the current?

hero (hero):

Hmmm, I could have sworn someone helped you out with this question before

hero (hero):

Use d = rt and r = (b ± c) to solve

hero (hero):

Upstream Equation: 248 = (b-c)4 Downstream Equation 368 = (b+c)4

hero (hero):

248/4 = b - c 368/4 = b + c 62 = b - c 92 = b + c b = 62 + c b = 92 - c 2b = 154 b = 154/2 b = 77 c = 15